OpenStudy (anonymous):

find the inverse of the matrix 3 -1 if it exists -10 9

OpenStudy (anonymous):

invese of \[\begin{matrix} a & b \\ c & d \end{matrix}\] is \[\begin{matrix} d & -b \\ -c & a \end{matrix}\] divided by the determinant let me see if i can write it better

OpenStudy (anonymous):

\[\frac{1}{ad-bc}\begin{pmatrix} d & -b \\ -c & a \end{pmatrix}\]

OpenStudy (anonymous):

that thing out front means divide by the determinant

OpenStudy (anonymous):

is it 9/17 1/17 ??? 10/17 3/17

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

hold on let me check

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

not too hard, was it?